OpenStudy (anonymous):

An expression is shown below: f(x) = -16x2 + 22x + 3 Part A: What are the x-intercepts of the graph of the f(x)? Show your work. Part B: Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work. Part C: What are the steps you would use to graph f(x)? Justify that you can use the answers obtained in Part A and Part B to draw the graph.

OpenStudy (anonymous):

PART A. set the function equal zero, and solve for x. -16x² + 22x + 3 = 0 16x² - 22x - 3 = 0 (8x + 1)(2x - 3)= 0 8x + 1= 0 8x=1 x= 1/8 2x - 3 =0 2x=3 x= 3/2

OpenStudy (anonymous):

I am too lazy/stupid (which ever one you pick) to do the rest

OpenStudy (anonymous):

You don't have to do the rest. lol Thanks for your help. (:

OpenStudy (anonymous):

You welcome :) Good luck !